OpenStudy (anonymous):

write a quadratic equation with the given roots in ax^2+bx+c=0 form (5/4) and 9

OpenStudy (cwrw238):

sum of roots = -b/a product of rots = c/a so here we get: 5/4 + 9 = -b/a so 41/4 = -b/a 45/4 = c/a

OpenStudy (anonymous):

or as an alternative method, having 5/4 and 9 as roots means the they are solutions to this equation: \(\large (x-\frac{5}{4})(x-9)=0 \) all you need to do is multiply it out....

OpenStudy (cwrw238):

yes - why didn't i think of that!!??

OpenStudy (anonymous):

lol well thanks for the help guys...but it doesn't really answer the question..i need to write an equation for specific roots....not to solve